In Okunland, a country whose economy operates according to Okun's law, real GDP equals $7,520 billion, potential GDP equals $8,000 billion, and the actual unemployment rate is 8 percent. What is the natural rate of unemployment in Okunland?

Multiple Choice

  • 11 percent

  • 2 percent

  • 5 percent

  • 6 percent
Answer #1

ANSWER:

Difference between potential and real output = 8,000 - 7,520 = 480

% of output below potential = (Difference between potential and real output / potential gdp) * 100 = (480 / 8000) * 100 = 0.06 * 100 = 6%

Each % above the natural rate of unemployment translates towards 2% negative output gap and , so the unemployment rate is 3% more then the natural rate.

natural rate = actual unemployment rate - 3% = 8% - 3% = 5%

so the correct answer is option c that is 5%.